A buffer was prepared by mixing 1.00 mole of ammonia and 1.00 mole of

ammonium chloride to form an aqueous solution with a total volume of 1.00

liter. To 500 mL of this solution was added 30.0 mL of 1.00 M NaOH. What

is the pH of this solution?


1
Expert's answer
2022-04-26T12:59:01-0400

Number of mol of NH3 = 500/1000 = 0.5mol

Number of moles of NH4Cl = 0.5 mol

Number of moles of NaOH = 30/1000 = 0.03mol

pKa of NH3 = 9.25 

pH = pKa + log (NH3/NH4+)

= 0.5/0.5 = 1 

Log 1 = 0

pH = 9.25 - 0 = 9.25 

∆ pH = 9.25 - 4.6 = 4.65